In the lungs, the respiratory membrane separates tiny sacs of air (pressure 1.00x105Pa) from the blood in 

the capillaries. These sacs are called alveoli. The average radius of the alveoli is 0.125 mm, and the air 

inside contains 14% oxygen. Assuming that the air behaves as an ideal gas at 310K, calculate the number 

of oxygen molecules in one of these sacs.


1
Expert's answer
2021-06-21T11:39:15-0400

"N_0=N\cdot0.14"


"pV=\frac{m}{M}RT\to m=\frac{pVM}{RT}=\frac{100000\cdot4/3\cdot3.14\cdot0.000125^3\cdot0.029}{8.31\cdot310}=9.2\cdot10^{-12}\ (kg)"


"m/M=N/N_A\to N=mN_A/M=9.2\cdot10^{-12}\cdot6.02\cdot10^{23}/0.029=1.91\cdot10^{14}"


"N_0=N\cdot0.14=1.91\cdot10^{14}\cdot0.14=2.7\cdot10^{13}" . Answer